OpenStudy (anonymous):

help please!! The table shows the population, p, of mice in a field at the end of m months. Time,m,:0,1,2,3,4 population,p,:6,12,24,48,96 Based on the data in the table, what will be the population of mice in the field at the end of 8 months? F 192 G 3,072 H 1,536 J 256

OpenStudy (anonymous):

its a geometric progression an = ar^(n-1) a= 6 r = 2 n = 8

OpenStudy (anonymous):

i got 12^7?

OpenStudy (anonymous):

and then find the sum \[\sum_{i=0}^{i=n} (ar ^{n}) = a((1-r ^{n})/(1-r))\]

OpenStudy (anonymous):

first find 2^ 7 then multiply with 6

OpenStudy (anonymous):

ok i got 768

OpenStudy (anonymous):

now find the sum by the given formula

OpenStudy (anonymous):

ugh now i have 6/7 I'm so sorry thanks for helping i just don't understand this unit

OpenStudy (anonymous):

simple way is use calculator and then start from 6 multiply it by 2 then multiply answer by 2 and so on till nine time. sum = 6(1-(2^8))/(1-2) = 6(-255)/(-1) = 6*255 = 1530

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

mention not